2.6 You move 3.27 m ahead, 2.00 m ahead, 7.95 backward 2.34 m ahead, 456 m backward, and 490 m ahcad. (a) What is your final displacement? (b) What is your maximum dis- placement? (c) What is your minimum displacement?
Answer #1

Let ahead be positive direction and backward be negative direction

move 1: 3.27 m ahead

Total displacement = 3.27 m

move 2: 2.00 m ahead

Total displacement = 3.27 + 2.00 = 5.27 m

move 3: 7.95 m backward

Total displacement = 5.27 - 7.95 = -2.68 m

move 4: 2.34 m ahead

Total displacement = -2.68 + 2.34 = -0.34 m

move 5: 4.56 m backward

Total displacement = -0.34 - 4.56 = -4.90 m

move 6: 4.90 m ahead

Total displacement = -4.90 + 4.90 = 0 m

a)

Final displacement = 0 m

Answer: 0

b)

Displacement is maximum after move 2 which is 5.27 m

Answer: 5.27 m

c)

Minimum displacement is 0 m

Answer: 0